Im taking algebra 1 and I don't understand at all Im stuck on this one problem I just can't solve it and I don't understand how to attempt to get the answer.
-5x + 3y = 11
x - 2y = 2
they want me to solve this problem using the substitution method?? can you please explain this to me
Thank you

Answer by rfer(16322)

x-2y=2 becomes x=2y+2
sutstitute 2y+2 for x in the other equation.
-5(2y+2)+3y=11
-10y-10+3y=11
combine
-7y=21
y=-3
---------------------
use -3 for y
-5x+3*-3=11
-5x-9=11
-5x=20
x=-4
----------------
put both values in the other equation for a check.
-4-2(-3)=2
-4+6=2